# AdapDISCOM: An Adaptive Sparse Regression Method for High-Dimensional Multimodal Data With Block-Wise Missingness and Measurement Errors

### v1

@@ 3.2 Adaptive direct sparse regression procedure using covariance from multimodality data (AdapDISCOM)
Note that, the optimal parameters of AdapDISCOM are a generalization of those of DISCOM and when $K=1,$ or when all parameters $\alpha_{k}^{*}$ are equal then Proposition 1 is the same as Proposition 1 shown in [54]. Furthermore, the expression of $\gamma^{*}$ shows how AdapDISCOM accounts for the heterogeneity of each modality’s errors, where the estimator of the intra-modality covariance matrix of modality $k$ can be considered to be $\alpha_{k}^{*}\widetilde{\Sigma}_{I_{k}}+w_{k}\operatorname{Tr}(\widetilde{\Sigma}_{I_{1}}/p)\mathbb{I}_{p},$ where $w_{k}=\frac{(1-\alpha_{k}^{*})^{2}}{\sum_{k=1}^{K}(1-\alpha_{k}^{*})^{2}}.$
@@
Thus, when a group of modalities is relatively less contaminated by measurement errors and missing data, then their intra-modality covariance matrix is relatively precise $(\delta_{I_{k}}^{2}\ \text{is small})$ and their optimal weight $\alpha_{k}^{*}=\frac{\theta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}$ should be large compared to other modalities, and the percentage of improvement is relatively small. This highlights the importance of using weights adapted to each modality, given that some modalities may be less contaminated than others and therefore have a more precise covariance. In the same way, the intra-modality covariance matrix is more accurate than the inter-modality covariance matrix, due to the unequal sample sizes [54].
@@
where $\psi_{H}(z)=z\cdot\mathds{1}_{\{|z|\leq H\}}+H\cdot\operatorname{sign}(z)\mathds{1}_{\{|z|>H\}}$ is the Huber function [24], $\mu$ the mean and $H$ an hyperparameter with default value chosen to be 1.345.1.345.1.345 . However, in this setting we propose to use flexibly different values of $H$ to account for the different numbers of samples available. The estimator built from the initial robust estimators will be designated AdapDISCOM-Huber in the following.
@@ 3.3.1 Sub-Gaussian Case
[…]
@@
[…]
@@
Condition A1 posits that both the predictors and the response variable follow sub-Gaussian distributions, while the RE condition A2 is used to derive statistical error bounds for the Lasso estimate [8]. The following theorem shows the large deviation bounds of $\widetilde{\Sigma}$ and $\widetilde{\boldsymbol{C}}$ as well as the convergence rate of $\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}.$
@@
$\widetilde{\text{C}}.$   Given the following constants $\nu_{3}=16(1+4\frac{\text{L}^{2}}{\min\{\operatorname{\mathbb{V}\text{ar}}(y),1\}})$ and $\nu_{4}=4$ we have
@@
With respect to the convergence of $\widetilde{\boldsymbol{\beta}},$ it can be shown that, under the assumptions of no missing data, Gaussian-distributed predictors, and independent and identically distributed Gaussian errors, we have from Theorem 1 that $\|\widetilde{\text{C}}-\widehat{\Sigma}\boldsymbol{\beta}^{0}\|_{\max}=\|\boldsymbol{X}{}^{\text{T}}\boldsymbol{\varepsilon}/n\|_{\max}=O_{p}(\sqrt{(\log p)/n)},$ and $\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}=O_{p}(\sqrt{(s\log p)/n)},$ which is the minimax $l_{2}$-norm rate as shown by Raskutti et al., [39]. Since the complete data generated from the Gaussian random design can be viewed as a special type of block-missing multimodality data, the error bound in Theorem 1 is sharp.
In general, the RE condition is satisfied when the design matrix X is fixed and complete. However, when this is not the case, for example, when the design matrix is Gaussian or random, additional assumptions are required to ensure the RE condition holds. For instance, in the case of a random design matrix, van de Geer and Bühlmann, [49] showed that the empirical covariance matrix $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ satisfies the RE condition as long as the true covariance matrix $\Sigma$ does, and $s^{2}\log p/n=o(1).$ Similarly, Raskutti et al., [39] demonstrated that, for a Gaussian random design matrix, $\widehat{\Sigma}$ satisfies the RE condition with high probability provided that the true Gaussian covariance matrix $\Sigma$ satisfies the RE condition and $n>\text{Constant}\cdot s\log p.$ Notably, their analysis is global, focusing on the full random matrix $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ rather than on individual entries of $\widehat{\Sigma}.$
In our work, as in DISCOM [54], we consider a general random design matrix encompassing both sub-Gaussian and heavy-tailed distributions, and we study a proposed estimated covariance matrix $\widehat{\Sigma}\neq\textbf{X}{}^{\text{T}}\textbf{X}/n$ in most cases. To ensure the RE condition holds with high probability, we adopt the assumption $\ s\sqrt{(\log p)/\min_{j,t,k}n_{jt}^{k}}=o(1),$ which closely parallels the condition $\ s^{2}\log p/n=o(1),$ used by van de Geer and Bühlmann, [49] for the complete data.
@@ 3.3.2 Heavy-tailed Case
(A3). Suppose that $\ \max_{1\leq j\leq p}\operatorname{\mathbb{E}}[\boldsymbol{x}_{j}^{4}]\leq Q_{1}^{2}/48$ and $\ \operatorname{\mathbb{E}}[\boldsymbol{y}^{4}]\leq Q_{2}^{2},$ where $Q_{1}$ and $Q_{2}$ are two positive constants.
In the Lasso literature, most studies assume a fixed design matrix [33, 57, 59] and Gaussian errors [55], or at most sub-exponential tails [6, 34]. In contrast, following the DISCOM approach, we consider a random design matrix and assume that the fourth moments of both the predictors and the response variable are finite. Note that under condition (A3) the tails of the distributions of $\boldsymbol{x}_{j}$’s and $\boldsymbol{y}$ may not be exponentially bounded.
@@
$\breve{\text{C}}.$   Furthermore, let $\text{H}_{j}^{k}=(\text{Q}_{1}+\text{Q}_{2})\sqrt{n_{j}^{k}/\log p}$ for each $j\in\{1,2,\ldots,p\},\ k\in\{1,2,\ldots,K\},$ we have
@@
Theorem 2 establishes that, under assumption (A3), the convergence rate of $\|\breve{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}$ in the heavy-tailed setting matches that of $\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}$ in the sub-Gaussian setting. Consistent with the findings in DISCOM [54], our Huber’s M-estimators in the heavy-tailed setting achieve the same convergence rate as the sample covariance-based estimators in the sub-Gaussian case. Furthermore, when the predictor and response distributions do not have exponentially bounded tails, the large deviation bounds for $\ \widetilde{\Sigma}$ and $\ \widetilde{C}$ can be wider than those of the Huber’s M-estimators $\ \breve{\Sigma}$ and $\ \breve{C},$ respectively. For additional details, we refer the reader to [54].
@@ 3.3.3 Model Selection Consistency
According to Yu et al., [54], condition (A4) can be viewed as the population-level counterpart of the strong irrepresentable condition proposed in Zhao and Yu, [57]. The following theorem show the model selection consistent for the AdapDISCOM method in the sub-Gaussian and heavy-tailed case.
@@
[…]
@@
[…]
@@
In the specific case of a Gaussian design, Wainwright, [50] showed that, by leveraging concentration inequalities for the normal distribution and the fact that $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ for the complete data, one can establish model selection consistency provided $n>\text{Constant}\cdot s\log(p-s).$ In our setting, however, we consider a general random design encompassing both sub-Gaussian and heavy-tailed distributions, and $\widehat{\Sigma}\neq\textbf{X}{}^{\text{T}}\textbf{X}/n$ due to the block-wise missingness in multimodal data. Therefore, we rely on $\|(\Sigma_{JJ})^{-1}\|_{\infty}\cdot\sqrt{\frac{s^{2}\log p}{\min_{j,t,k}n_{jt}^{k}}}=o(1)$ to ensure that properties $\|(\widehat{\Sigma}_{JJ})^{-1}\|_{\infty}\leq\text{Constant}\cdot\|(\Sigma_{JJ})^{-1}\|_{\infty}$ and $\|\widehat{\Sigma}_{J^{c}J}\widehat{\Sigma}_{JJ}^{-1}\|_{\infty}\leq 1-\eta\prime$ are satisfied with high probability, if $\|\Sigma_{J^{c}J}\Sigma_{JJ}^{-1}\|_{\infty}\leq 1-\eta.$ The parameters $\eta\prime\in(0,1)$ and $\eta\in(0,1)$ are two constants. It is worth noting that this condition has also been used to establish model selection consistency for random designs, as in [26].
@@ 3.3.4 Efficient estimation in the Optimal Linear Prediction
[…]
@@
we have $l_{\min}$ always less than $l_{\max}=1/m_{C}.$ Therefore, instead of having to estimate $K+2$ hyperparameters, this parametrization reduces the tuning to the search for the best $l_{0}\in[l_{\min},l_{\max}]$ and the parameter $\lambda.$ Furthermore, rather than performing the eigendecomposition for each parameter combination to ensure that $\widehat{\Sigma}$ is positive semidefinite, this fast reparametrization method requires two eigendecomposition of $\widetilde{\Sigma}$ and the matrix from (8) before the tuning parameter selection process. Similar to [54], for each $l_{0}\in[l_{\min},l_{\max}],$ we can incorporate the coordinate descent algorithm [18] on a grid of $\lambda$ values or apply the LARS algorithm [26] to compute the solution path.
The resulting method from this procedure is referred to as FastAdapDISCOM. Note that, when $k=1$ or $m_{k}=m$ for all $k\in\{1,2,\ldots,K\}$ then FastAdapDISCOM reduces to the FastDISCOM method proposed by [54].
@@ 4.1 Design
Following the multimodal framework proposed by Yu et al., [54], we assume that X is a block-missing matrix comprising three modalities. Specifically, X consists of $p=300$ predictors, evenly distributed across the three modalities (100100100 predictors per modality). The training sample size is $n,$ with $n/4$ missing values in each modality. That is, the training dataset is composed of $n/4$ samples with complete observations, $n/4$ samples with observations from the first and the second modalities, $n/4$ samples with observations from the first and the third modalities, and $n/4$ samples with missing observations only from the first modality. See the supplementary material for an illustration of the missingness pattern. We consider different sample sizes $n\in\{40,120,200,280,360,440,520\},$ allowing us to evaluate model performance in both high-dimensional $(n<p)$ and standard $(n\geq p)$ settings. The validation and test sets consist of 200200200 and 400400400 samples with complete observations, respectively. The coefficient vector is $\boldsymbol{\beta}^{0},$ with each modality containing five nonzero predictors.
@@
We consider six distinct block-missing scenarios to analyze the effect of multimodal heterogeneous data structure and heterogeneous measurement error contamination across modalities. In Scenario I, the modalities are homogeneous and subject to the same level of measurement error. Specifically, the three block-missing modalities are generated by the same multivariate Gaussian distribution $\mathcal{N}(0,\Sigma=(\sigma_{jt}^{2})_{1\leq j,t\leq p}),$ where $\sigma_{jt}^{2}=0.6^{|j-t|}.$ The additive measurement error is identical across modalities and takes values $\tau^{2}\in\{0,0.2,0.4,0.6,0.8\}.$ This scenario is designed to isolate the impact of missing data and measurement error in the absence of structural heterogeneity across modalities. Note that when $\tau^{2}=0,$ the data are only subject to missingness, corresponding to the design considered by Yu et al., [54].
@@
  • Modality 2 features a block-diagonal correlation structure, constructed by repeating $p/5$ times the $\times 5$ 5 block matrix $\Sigma_{\text{bloc}}=0.15\cdot\mathbf{1}_{5\times 5}+0.85\cdot\mathbb{I}_{5},$ as in Example 2 of the simulation design in Yu et al., [54]
  • Modality 3, inspired by Li et al., [28], is generated from a covariance matrix defined as the Kronecker product of two matrices: $\Sigma=\Sigma_{1}\otimes\Sigma_{2}$, where $\Sigma_{1}$ is $\times 3$ 3 matrix with entries $\Sigma_{1jt}=0.8^{|j-t|},$ and $\Sigma_{2}$ is a larger $p/3\times p/3$ matrix with entries $\Sigma_{2jt}=0.3^{|j-t|}.$
@@
Each simulation is repeated $B=100$ times, and variability in the results is visualized using boxplots. We compare the performance of AdapDISCOM-based methods against nine competing approaches: the DISCOM-based methods (DISCOM, Fast-DISCOM, DISCOM-Huber, and Fast-DISCOM-Huber), SCOM [28], CocoLasso [8], two LASSO-based imputation methods (LASSO-mean and LASSO-SVD), and the complete-case LASSO [45], which uses only complete observed samples. Performance is evaluated using four metrics:
@@
  • Bias, computed as $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2},$ which quantifies the estimation error between the true parameter vector and its estimate [21].
  • F1-score, defined as $F_{1}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}},$ which evaluates the variable selection accuracy, where $\text{TP},\ FP,\ \text{and}\ FN$ denote the numbers of true positives, false positives, and false negatives, respectively [43].

### v2

@@ 3.2 Adaptive direct sparse regression procedure using covariance from multimodality data (AdapDISCOM)
Note that, the optimal parameters of AdapDISCOM are a generalization of those of DISCOM and when $K=1,$ or when all parameters $\alpha_{k}^{*}$ are equal then Proposition 1 is the same as Proposition 1 shown in [Yu et al., 2020]. Furthermore, the expression of $\gamma^{*}$ shows how AdapDISCOM accounts for the heterogeneity of each modality’s errors, where the estimator of the intra-modality covariance matrix of modality $k$ can be considered to be $\alpha_{k}^{*}\widetilde{\Sigma}_{I_{k}}+w_{k}\operatorname{Tr}(\widetilde{\Sigma}_{I_{1}}/p)\mathbb{I}_{p},$ where $w_{k}=\frac{(1-\alpha_{k}^{*})^{2}}{\sum_{k=1}^{K}(1-\alpha_{k}^{*})^{2}}.$
@@
Thus, when a group of modalities is relatively less contaminated by measurement errors and missing data, then their intra-modality covariance matrix is relatively precise $(\delta_{I_{k}}^{2}\ \text{is small})$ and their optimal weight $\alpha_{k}^{*}=\frac{\theta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}$ should be large compared to other modalities, and the percentage of improvement is relatively small. This highlights the importance of using weights adapted to each modality, given that some modalities may be less contaminated than others and therefore have a more precise covariance. In the same way, the intra-modality covariance matrix is more accurate than the inter-modality covariance matrix, due to the unequal sample sizes [Yu et al., 2020].
@@
where $\psi_{H}(z)=z\cdot\mathds{1}_{\{|z|\leq H\}}+H\cdot\operatorname{sign}(z)\mathds{1}_{\{|z|>H\}}$ is the Huber function [Huber, 1964], $\mu$ the mean and $H$ an hyperparameter with default value chosen to be $1.345.$ However, in this setting we propose to use flexibly different values of $H$ to account for the different numbers of samples available. The estimator built from the initial robust estimators will be designated AdapDISCOM-Huber in the following.
@@ 3.3.1 Sub-Gaussian Case
$(\textbf{A1}).$ Suppose that there exists a constant $L>0$ such that
@@
$(\textbf{A2}).$ Suppose that the true covariance matrix $\Sigma$ satisfies the following restricted eigenvalue (RE) condition:
@@
Condition A1 posits that both the predictors and the response variable follow sub-Gaussian distributions, while the RE condition A2 is used to derive statistical error bounds for the Lasso estimate [Datta and Zou, 2017]. The following theorem shows the large deviation bounds of $\widetilde{\Sigma}$ and $\widetilde{\boldsymbol{C}}$ as well as the convergence rate of $\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}.$
@@
[…]
@@
With respect to the convergence of $\widetilde{\boldsymbol{\beta}}$, it can be shown that, under the assumptions of no missing data, Gaussian-distributed predictors, and independent and identically distributed Gaussian errors, we have from Theorem 1 that $\|\widetilde{\text{C}}-\widehat{\Sigma}\boldsymbol{\beta}^{0}\|_{\max}=\|\boldsymbol{X}{}^{\text{T}}\boldsymbol{\varepsilon}/n\|_{\max}=O_{p}(\sqrt{(\log p)/n)},$ and $\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2}=O_{p}(\sqrt{(s\log p)/n)},$ which is the minimax $l_{2}$-norm rate as shown by Raskutti et al. [2011]. Since the complete data generated from the Gaussian random design can be viewed as a special type of block-missing multimodality data, the error bound in Theorem 1 is sharp.
In general, the RE condition is satisfied when the design matrix X is fixed and complete. However, when this is not the case, for example, when the design matrix is Gaussian or random, additional assumptions are required to ensure the RE condition holds. For instance, in the case of a random design matrix, van de Geer and Bühlmann [2009] showed that the empirical covariance matrix $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ satisfies the RE condition as long as the true covariance matrix $\Sigma$ does, and $s^{2}\log p/n=o(1).$ Similarly, Raskutti et al. [2011] demonstrated that, for a Gaussian random design matrix, $\widehat{\Sigma}$ satisfies the RE condition with high probability provided that the true Gaussian covariance matrix $\Sigma$ satisfies the RE condition and $n>\text{Constant}\cdot s\log p.$ Notably, their analysis is global, focusing on the full random matrix $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ rather than on individual entries of $\widehat{\Sigma}$.
In our work, as in DISCOM [Yu et al., 2020], we consider a general random design matrix encompassing both sub-Gaussian and heavy-tailed distributions, and we study a proposed estimated covariance matrix $\widehat{\Sigma}\neq\textbf{X}{}^{\text{T}}\textbf{X}/n$ in most cases. To ensure the RE condition holds with high probability, we adopt the assumption $s\sqrt{(\log p)/\min_{j,t,k}n_{jt}^{k}}=o(1),\$which closely parallels the condition $s^{2}\log p/n=o(1),\$used by van de Geer and Bühlmann [2009] for the complete data.
@@ 3.3.2 Heavy-tailed Case
(A3). Suppose that $\max_{1\leq j\leq p}\operatorname{\mathbb{E}}[\boldsymbol{x}_{j}^{4}]\leq Q_{1}^{2}/48\$and $\operatorname{\mathbb{E}}[\boldsymbol{y}^{4}]\leq Q_{2}^{2},\$where $Q_{1}$ and $Q_{2}$ are two positive constants.
In the Lasso literature, most studies assume a fixed design matrix [Meinshausen and Bühlmann, 2006; Zhao and Yu, 2006; Zou, 2006] and Gaussian errors [Zhang and Huang, 2008], or at most sub-exponential tails [Bunea, 2008; Meinshausen and Yu, 2009]. In contrast, following the DISCOM approach, we consider a random design matrix and assume that the fourth moments of both the predictors and the response variable are finite. Note that under condition (A3) the tails of the distributions of $\boldsymbol{x}_{j}$’s and $\boldsymbol{y}$ may not be exponentially bounded.
@@
[…]
@@
Similarly, when $p$ is fixed, comparable bounds can be obtained in the sub-Gaussian setting. Moreover, as shown by Yu et al. [2020], the convergence rate of the estimation error in the classical fixed-$p$ regime is significantly faster than in the high-dimensional setting where $p\longrightarrow\infty.$
@@ 3.3.3 Model Selection Consistency
According to Yu et al. [2020], condition (A4) can be viewed as the population-level counterpart of the strong irrepresentable condition proposed in Zhao and Yu [2006]. The following theorem show the model selection consistent for the AdapDISCOM method in the sub-Gaussian and heavy-tailed case.
@@
Sub-Gaussian Case.   Under conditions $(\textbf{A1})$ and $(\textbf{A4}),$ let $1-\alpha_{k}=O\Big(\sqrt{(\log p/\min_{j}n_{j}^{k})}\Big)$ and $1-\alpha_{C}=O\Big(\sqrt{(\log p/\min_{j,t,k}n_{jt}^{k})}\Big),$   If
@@
Heavy-Tailed Case.   Under conditions $(\textbf{A3})$ and $(\textbf{A4}),$ let $1-\alpha_{k}=O\Big(\sqrt{(\log p/\min_{j}n_{j}^{k})}\Big)$ and $1-\alpha_{C}=O\Big(\sqrt{(\log p/\min_{j,t,k}n_{jt}^{k})}\Big),\;\text{H}_{jt}^{k}=\frac{\text{Q}_{1}}{12}\sqrt{n_{jt}^{k}/\log p}$ and   $\text{H}_{j}^{k}=(\text{Q}_{1}+\text{Q}_{2})\sqrt{n_{j}^{k}/\log p}.$   If
@@
In the specific case of a Gaussian design, Wainwright [2009] showed that, by leveraging concentration inequalities for the normal distribution and the fact that $\widehat{\Sigma}=\textbf{X}{}^{\text{T}}\textbf{X}/n$ for the complete data, one can establish model selection consistency provided $n>\text{Constant}\cdot s\log(p-s).$ In our setting, however, we consider a general random design encompassing both sub-Gaussian and heavy-tailed distributions, and $\widehat{\Sigma}\neq\textbf{X}{}^{\text{T}}\textbf{X}/n$ due to the block-wise missingness in multimodal data. Therefore, we rely on $\|(\Sigma_{JJ})^{-1}\|_{\infty}\cdot\sqrt{\frac{s^{2}\log p}{\min_{j,t,k}n_{jt}^{k}}}=o(1)$ to ensure that properties $\|(\widehat{\Sigma}_{JJ})^{-1}\|_{\infty}\leq\text{Constant}\cdot\|(\Sigma_{JJ})^{-1}\|_{\infty}$ and $\|\widehat{\Sigma}_{J^{c}J}\widehat{\Sigma}_{JJ}^{-1}\|_{\infty}\leq 1-\eta\prime$ are satisfied with high probability, if $\|\Sigma_{J^{c}J}\Sigma_{JJ}^{-1}\|_{\infty}\leq 1-\eta.$ The parameters $\eta\prime\in(0,1)$ and $\eta\in(0,1)$ are two constants. It is worth noting that this condition has also been used to establish model selection consistency for random designs, as in [Jeng and Daye, 2011].
@@ 3.3.4 Efficient estimation in the Optimal Linear Prediction
Given initial estimates of $\Sigma$ and $\textbf{C},$ for example $\widetilde{\Sigma}=\sum_{k=1}^{K}\widetilde{\Sigma}_{I_{k}}+\widetilde{\Sigma}_{C}$ and $\widetilde{\textbf{C}}$, the proposed AdapDISCOM method estimates $\boldsymbol{\beta}^{0}$ by solving the following optimization problem:
@@
we have $l_{\min}$ always less than $l_{\max}=1/m_{C}.$ Therefore, instead of having to estimate $K+2$ hyperparameters, this parametrization reduces the tuning to the search for the best $l_{0}\in[l_{\min},l_{\max}]$ and the parameter $\lambda.$ Furthermore, rather than performing the eigendecomposition for each parameter combination to ensure that $\widehat{\Sigma}$ is positive semidefinite, this fast reparametrization method requires two eigendecomposition of $\widetilde{\Sigma}$ and the matrix from (8) before the tuning parameter selection process. Similar to [Yu et al., 2020], for each $l_{0}\in[l_{\min},l_{\max}],$ we can incorporate the coordinate descent algorithm [Friedman et al., 2010] on a grid of $\lambda$ values or apply the LARS algorithm [Jeng and Daye, 2011] to compute the solution path.
The resulting method from this procedure is referred to as FastAdapDISCOM. Note that, when $k=1$ or $m_{k}=m$ for all $k\in\{1,2,\ldots,K\}$ then FastAdapDISCOM reduces to the FastDISCOM method proposed by [Yu et al., 2020].
@@ 4.1 Design
Following the multimodal framework proposed by Yu et al. [2020], we assume that X is a block-missing matrix comprising three modalities. Specifically, X consists of $p=300$ predictors, evenly distributed across the three modalities ($100$ predictors per modality). The training sample size is $n,$ with $n/4$ missing values in each modality. That is, the training dataset is composed of $n/4$ samples with complete observations, $n/4$ samples with observations from the first and the second modalities, $n/4$ samples with observations from the first and the third modalities, and $n/4$ samples with missing observations only from the first modality. See the supplementary material for an illustration of the missingness pattern. We consider different sample sizes $n\in\{40,120,200,280,360,440,520\},$ allowing us to evaluate model performance in both high-dimensional $(n<p)$ and standard $(n\geq p)$ settings. The validation and test sets consist of $200$ and $400$ samples with complete observations, respectively. The coefficient vector is $\boldsymbol{\beta}^{0},$ with each modality containing five nonzero predictors.
@@
We consider six distinct block-missing scenarios to analyze the effect of multimodal heterogeneous data structure and heterogeneous measurement error contamination across modalities. In Scenario I, the modalities are homogeneous and subject to the same level of measurement error. Specifically, the three block-missing modalities are generated by the same multivariate Gaussian distribution $\mathcal{N}(0,\Sigma=(\sigma_{jt}^{2})_{1\leq j,t\leq p}),$ where $\sigma_{jt}^{2}=0.6^{|j-t|}.$ The additive measurement error is identical across modalities and takes values $\tau^{2}\in\{0,0.2,0.4,0.6,0.8\}.$ This scenario is designed to isolate the impact of missing data and measurement error in the absence of structural heterogeneity across modalities. Note that when $\tau^{2}=0,$ the data are only subject to missingness, corresponding to the design considered by Yu et al. [2020].
@@
  • Modality 2 features a block-diagonal correlation structure, constructed by repeating $p/5$ times the $5\times 5$ block matrix $\Sigma_{\text{bloc}}=0.15\cdot\mathbf{1}_{5\times 5}+0.85\cdot\mathbb{I}_{5},$ as in Example 2 of the simulation design in Yu et al. [2020]
  • Modality 3, inspired by Li et al. [2023], is generated from a covariance matrix defined as the Kronecker product of two matrices: $\Sigma=\Sigma_{1}\otimes\Sigma_{2}$, where $\Sigma_{1}$ is $3\times 3$ matrix with entries $\Sigma_{1jt}=0.8^{|j-t|},$ and $\Sigma_{2}$ is a larger $p/3\times p/3$ matrix with entries $\Sigma_{2jt}=0.3^{|j-t|}.$
@@
Each simulation is repeated $B=100$ times, and variability in the results is visualized using boxplots. We compare the performance of AdapDISCOM-based methods against nine competing approaches: the DISCOM-based methods (DISCOM, Fast-DISCOM, DISCOM-Huber, and Fast-DISCOM-Huber), SCOM [Li et al., 2023], CocoLasso [Datta and Zou, 2017], two LASSO-based imputation methods (LASSO-mean and LASSO-SVD), and the complete-case LASSO [Tibshirani, 1996], which uses only complete observed samples. Performance is evaluated using four metrics:
@@
  • Bias, computed as $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{0}\|_{2},$ which quantifies the estimation error between the true parameter vector and its estimate [Hastie et al., 2009].
  • F1-score, defined as $F_{1}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}},$ which evaluates the variable selection accuracy, where $\text{TP},\ FP,\ \text{and}\ FN$ denote the numbers of true positives, false positives, and false negatives, respectively [Taha and Hanbury, 2015].